OpenStudy (anonymous):

When comparing twenty-three forty-seconds to one-half using mental math, what fraction is useful? twenty-three halves twenty-three twenty-firsts twenty-one forty-seconds twenty-three eighty-forths

OpenStudy (anonymous):

In forty-seconds how would you express 1/2? That is \[\frac{ 1 }{ 2 }=\frac{ ? }{ 42 }\]